OpenStudy (anonymous):

1. On a vertical wall a certain painting, whose base is 6 ft. from the floor, is hanged. From a point on the floor, the angles of elevation of the top and bottom of the flagpole are 45° and 30 ° . Find the vertical dimension of the painting. 2. The angle of elevation from the top of the cliff to the top of the second cliff 290 ft. away is 68° , while the angle of depression from the top of the the first cliff to the bottom of the second cliff is 63° . Find the height of the second cliff.

OpenStudy (anonymous):

could someone give me the whole picture about that problem and probably I can solve it.huhuhu

OpenStudy (anonymous):

the answer of that problem is 4.4 ft but i cannot get the right solution.huhuhu.

OpenStudy (anonymous):

it's our homework and our prof want us to answer that.

OpenStudy (zpupster):

 angle of elevation problem lets find the point on floor from this tan x= height/ distance to point on floor tan (30 degrees) = 6 feet/ pof - rearrange pof = 6/tan 30 pof = 10.4 now we know the point on floor tan(45) = height/ 10.4 (1)(10.4) = height on wall lets subtract 10.4- 6 = 4.4 angle of depression problems are the same way

OpenStudy (anonymous):

omg. thank you very much.

OpenStudy (zpupster):

 (290) /(tan 68)= height on top portion angle of depression: tan= ht of cliff/distance tan(63)=h/290 tan63*(290)=height of bottom portion ht top+ht bottom = total ht of cliff

OpenStudy (zpupster):

did you get this i see you asked another here: http://openstudy.com/users/lynndale#/updates/5129a526e4b098bb5fba7675

OpenStudy (anonymous):

Yap i was able to answer the cliff problem 1286.94 ft. By the way, thanks for the solution. :D